OpenStudy (anonymous):

Victor calculates the percentage interest rate he needs to make twenty dollars in simple annual interest over 2 months on a deposit of one thousand dollars. Where did he go wrong? I equals p times r times t; twenty equals one thousand times r times two; r equals zero point zero one; r equals one percent He applied the principal and interest to the wrong variables. He should have divided 2 by twelve to get the value for t. He misrepresented the earned interest with the final balance. He should have multiplied r by one hundred to get the interest rate as a percentage.

OpenStudy (anonymous):

do we have to do calculation and working here or it's just a MCQ?

OpenStudy (anonymous):

No I don't think so. It's a multiple choice question. What is Victor's error? I need the right answer.

OpenStudy (anonymous):

FIrst of all ther's a typo: \[I != P \times R \times T\] \[I = \frac{ P \times R \times T }{ 100 }\]

OpenStudy (anonymous):

There is?

OpenStudy (anonymous):

don't believe it then google it

OpenStudy (anonymous):

No I believe you so then what is the answer if there is a typo?

OpenStudy (anonymous):

He should have divided 2 by 12

OpenStudy (anonymous):

2nd option

OpenStudy (anonymous):

Thank you so much! :) May you help with a little more? :S

OpenStudy (anonymous):

but I personalyy believe \[20 != \frac{ 1000 \times 1 \times 2 }{ 100 * 12 }= 10 /6 = 5/ 3\] @_lilb_to_smooth_4_U_ :that's no possible

OpenStudy (anonymous):

Oh... Yea I believe you. And also can you help me with a little more?

OpenStudy (anonymous):

with what?

OpenStudy (anonymous):

Same subject Math. Multiple choice questions.

OpenStudy (anonymous):

ok I 'll try ..post it

OpenStudy (anonymous):

Fayed tipped the pizza delivery driver three dollars for pizzas that cost eighteen dollars. Approximately what percent did he tip the driver? Round your answer to the nearest percent. three percent six percent fifteen percent seventeen percent

OpenStudy (anonymous):

very simple question: I ssume that you mean the percent is approcimated on the basis of the CP of pizza(yummy:P) Therefore: \[= \frac{ 3 }{ 18 } \times 100 = \frac{ 100 }{ 6} = \frac{ 50 }{ 3 } = 16.6 \approx 17\] Ans 17%

OpenStudy (anonymous):

Oh thank you so much!

OpenStudy (anonymous):

don't mention bro

OpenStudy (anonymous):

May you help me with more?

OpenStudy (anonymous):

as I already said I'll try ..:K post them